### v1

@@ 8.2 Parallel scalability
To evaluate the parallel performance of the code, we use the 120m-spaced grid and measure strong scaling from 64 to 2,048 processors. This mesh consists of a periodic tidal boundary and forcing from Hurricane Ike. A simulation duration of 1 day and $\Delta t=1$ s were used. We run both CG and DG-CG; for CG we use the explicit solver option and therefore no iterative matrix solve is performed. In Table 6, we show the runtime breakdown and speedup. This data is also plotted in Figure 19. We use the Tuning and Analysis Utilities (TAU) to measure the time. Most of the communication takes place during neighbor exchange of updated variables, and scale quite well with more processors. However, the overhead is still significant compared to the CG code because we now update additional variables like wet/dry flags and elevation before/after slope limiting.

### v2

@@ 7.2 Parallel scalability
To evaluate the parallel performance of the code, we use the 120m-spaced grid and measure strong scaling from 64 to 2,048 processors. This mesh consists of a periodic tidal boundary and forcing from Hurricane Ike. A simulation duration of 1 day and $\Delta t=1$ s were used. We run both CG and DG-CG; for CG we use the explicit solver option and therefore no iterative matrix solve is performed. In Table 6, we show the runtime breakdown and speedup. This data is also plotted in Figure 18. We use the Tuning and Analysis Utilities (TAU) [Shende2006-mq] to measure the time. Most of the communication takes place during neighbor exchange of updated variables, and scale quite well with more processors. However, the overhead is still significant compared to the CG code because we now update additional variables like wet/dry flags and elevation before/after slope limiting.